A scalloped hammerhead shark swims at a steady speed of 1.4 m/s with its 85-cm-wide head perpendicular to the earth's 60 μT magnetic field. What is the magnitude of the emf induced between the two sides of the shark's head? Answer in mV.
Answer #1

V = 1.4m/s , L = 0.85m , B = 60×10^-6T

Emf = change in flux/change in time

=) Emf = BVL = 60×10^-6 × 1.4 × 0.85 = 71.4×10^-6V

=) Emf = 0.0714 mV